OpenStudy (anonymous):

having trouble understanding how to sketch a plane curve represented by a vector valued function. ex: r(t)=t^2i+tj

OpenStudy (amistre64):

r is the distance from the origin; as defined by the i and j components

OpenStudy (amistre64):

or simply put r = <t^2,t>

OpenStudy (amistre64):

x = t^2 y = t +- sqrt(x) = t y = +-sqrt(x)

OpenStudy (anonymous):

so you always just convert it into x and y coordinates and graph it like normal?

OpenStudy (amistre64):

r(0) = (0,0) r(1) = (1,1) r(2) = (4,2) r(3) = (9,3)

OpenStudy (amistre64):

well, yes and no :) depends on the relation between x and y

OpenStudy (amistre64):

in this case; x is a function of t and y is a function of t so there is some relation between the 2

OpenStudy (amistre64):

and it looks like it creates: y = sqrt(x)

OpenStudy (anonymous):

so then should i be able to do this as long as x and y are both functions of t?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok, well it looks like that is all i will deal with for now so i guess that makes better sense. i just get caught off gaurd when i see another variable.

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

your welcome